OpenStudy (lilai3):

PLEASE HELP!! WILL REWARD MEDAL/FAN!!! Mature soil is composed of three layers, the uppermost being topsoil. Jamal is planting a bush that needs a hole 18 centimeters deep for the roots. The instructions suggest an additional 8 centimeters depth for a cushion. If Jamal wants to add even more cushion, and the topsoil in his yard is 30 centimeters deep, how much more cushion can he add and still remain in the topsoil layer? Thank you!~

OpenStudy (jjuden):

This is simple subtraction use your head. Topsoil = 30 cm Hole = 18 cm 30 - 18 = 12 cm for cushion Suggested 8 cm of cushion 12 - 8 = 4 cm Jamal can add 4 extra cm of cushion while remaining in the topsoil layer. 30 - 18 = 12 12-8= 4 he can add 4 more centimeters of cushion if he wants to remain in the topsoil layer.

OpenStudy (lilai3):

is it 18+x>30?

OpenStudy (anonymous):

Yes^^

OpenStudy (jjuden):

yes it is

OpenStudy (lilai3):

thank you so much! i would love to give both of you medals, but i think @jjuden should get it because of the explaining. thanks to both of you, though!

OpenStudy (jjuden):

i gave @its_ashley_doee one for you

OpenStudy (lilai3):

OMG thank you! ur sweet!

OpenStudy (anonymous):

Lol thanks guys! :)

OpenStudy (jjuden):

UR WELCOME :3